OpenStudy (anonymous):

Find the measure of each interior angle of a regular pentagon. answers - 176°, 108°, 120°, 150°

OpenStudy (anonymous):

Sum of all interior angles=(n-2)180 where n=no. of sides For regular pentagon, sum of all interior angles is 540 so one angle=540/5=108

OpenStudy (anonymous):

sum of exterior angles=360 each exterior angle of a regular polygon of n sides=360/n